# Constraining the Cosmic-ray Energy Based on Observations of Nearby Galaxy Clusters by LHAASO

### v1

@@ Abstract
Galaxy clusters act as reservoirs of high-energy cosmic rays (CRs). As CRs propagate through the intracluster medium, they generate diffuse $\gamma$-rays detectable by arrays such as LHAASO. These $\gamma$-rays result from proton-proton ($pp$) collisions of very high-energy cosmic rays (VHECRs) or inverse Compton (IC) scattering of positron-electron pairs created by $p\gamma$ interactions of ultra-high-energy cosmic rays (UHECRs). We analyzed diffuse $\gamma$-ray emission from the Coma, Perseus, and Virgo clusters using LHAASO data. Diffuse emission was modeled as a disk of radius $R_{500}$ for each cluster while accounting for point sources. No significant diffuse emission was detected, yielding 95% confidence level (C.L.) upper limits on the $\gamma$-ray flux: for WCDA (1–25 TeV) and KM2A ($>$25 TeV), less than (49.4, 13.7, 54.0) and (1.34, 1.14, 0.40) ×$\times 10^{-14}$ ph cm-2 s-1 for Coma, Perseus, and Virgo, respectively. The $\gamma$-ray upper limits can be used to derive model-independent constraints on the integral energy of CRp above 10 TeV (corresponding to the LHAASO observational range $>1\,\mathrm{TeV}$ under the $pp$ scenario) to be less than $(1.96,0.59,0.08)\times 10^{61}\,\mathrm{erg}$. The absence of detectable annuli/ring-like structures, indicative of cluster accretion or merging shocks, imposes further constraints on models in which the UHECRs are accelerated in the merging shocks of galaxy clusters.
@@ 1 Introduction
Galaxy clusters, recognized as the largest gravitationally bound entities in the universe, serve as vital cosmological laboratories for investigating a wide range of astronomical phenomena. These clusters typically encompass hundreds to thousands of galaxies, all enveloped in hot, diffuse intracluster gas that extends across several megaparsecs. Radio observations have uncovered non-thermal radio halos and relics within these clusters, indicating the presence of a population of relativistic particles (Feretti et al., 2012).
@@
Cosmic rays within galaxy clusters are generated and accelerated by multiple astrophysical sources. These include supernova-driven galactic winds, active galactic nuclei (AGNs) (Hinton et al., 2007), large-scale intergalactic shocks arising from accretion and merger processes (Colafrancesco & Blasi, 1998; Ryu et al., 2003; Vannoni et al., 2011; Brunetti & Jones, 2014), and turbulent intracluster magnetic fields (Brunetti & Blasi, 2005). The magnetic fields in galaxy clusters, typically on the order of a few $\mu G$, effectively confine cosmic rays, allowing their accumulation (Völk et al., 1996; Berezinsky et al., 1997; Govoni & Feretti, 2004). When these relativistic particles interact with the ambient matter, they can generate detectable high-energy $\gamma$-rays.
@@
Studying galaxy clusters in the $\gamma$-ray band provides crucial insights into the efficiency of particle acceleration, magnetic confinement, and radiation processes involving cosmic rays. Several radiation mechanisms can give rise to $\gamma$-rays in galaxy clusters. The hadronic scenario, which involves inelastic collisions between cosmic ray protons (CRp) and thermal nuclei (Dennison, 1980), is a major potential $\gamma$-ray radiation mechanism in galaxy clusters. Leptonic scenarios can also generate $\gamma$-rays. Relativistic CRp interact with target photon fields, such as the cosmic microwave background (CMB) and infrared background, producing high-energy $\gamma$-rays (Atoyan & Völk, 2000). $\gamma$-ray emission can also arise from IC scattering of secondary electrons, which are produced by Bethe-Heitler processes involving relativistic protons and the CMB photon field, if UHECRs are accelerated within galaxy clusters, as predicted by some models (Kelner & Aharonian, 2008). Another mechanism involves IC radiation emitted by relativistic cosmic ray electrons (CRe), which originate from interactions between very high-energy (VHE) $\gamma$-rays and the diffuse extragalactic background radiation field.
@@
Previous observations of $\gamma$-rays in the MeV to GeV energy range by EGRET (Reimer et al., 2003) provided only upper limits for several galaxy clusters. Similarly, the Whipple telescope reported upper limits at TeV energies for the Perseus and Abell 2029 clusters (Perkins et al., 2006). More recently, stereoscopic instruments such as H.E.S.S. and VERITAS have also set upper limits from TeV observations (HESS Collaboration et al., 2023; HESS Collaboration & Aharonian, 2009; Arlen et al., 2012; Perkins et al., 2006). The analysis of Fermi LAT data by Xi et al. (2018) revealed extended GeV emission from the direction of the Coma cluster, which was later confirmed by Baghmanyan et al. (2022).
@@
In this Letter, we report the LHAASO observations of three nearby galaxy clusters: Coma, Perseus, and Virgo. Leveraging the $\gamma$-ray upper limits obtained from LHAASO observations, we impose stringent constraints on the energy budget of high-energy particles within these clusters above 10 TeV, corresponding to the $\gamma$-ray energy range observed by LHAASO ($>1\,\mathrm{TeV}$) under the $pp$ interaction scenario, in a model-independent manner.
@@
The structure of this paper is as follows. In Sec.2, we present the fundamental background information on the selected clusters. In Sec.3, we conduct an in-depth analysis of the LHAASO-WCDA and LHAASO-KM2A data. Finally, in Sec.4, we discuss the derived cosmic-ray constraints and their astrophysical implications.
2 Facts of the galaxy clusters
@@
Coma, Perseus, and Virgo are the three closest massive galaxy clusters, with redshifts of approximately 0.023 (Giovannini et al., 1993), 0.017 (Hitomi Collaboration et al., 2018), and 0.004 (Mei et al., 2007), and total masses $M_{500}$ of about $6.13\times 10^{14}M_{\odot}$ (Planck Collaboration et al., 2013), $5.77\times 10^{14}M_{\odot}$ (CTA Consortium et al., 2023), and $0.83\times 10^{14}M_{\odot}$ (Simionescu et al., 2017), respectively (as shown in Tab.A.1). Their proximity and large mass render these clusters ideal targets for studying cosmic rays and dark matter.
@@
Observations in other wavelengths have revealed that all three clusters exhibit radio halos, which are a clear indication of a substantial population of relativistic particles. Specifically, the Coma cluster possesses a giant radio halo, which is typically associated with cluster mergers (Cassano et al., 2010). This halo extends from the core of the Coma cluster and features diffuse radio relics near its virial radius (Thierbach et al., 2003). X-ray observations have shown that NGC 4839 is a subgroup in the south-west region of the Coma cluster, connected to its core by a faint X-ray bridge, suggesting a recent merger event (Churazov et al., 2021; Malavasi et al., 2020). The presence of the giant radio halo, in tandem with the merger activities, implies that shock wave dissipation and turbulence resulting from the merging process play significant roles in injecting non-thermal energy into the system (Vannoni et al., 2011).
@@
Perseus and Virgo are well-known as relaxed cool-core (CC) clusters (Peterson & Fabian, 2006), characterized by their dense cores. Perseus showcases a radio mini-halo (RMH) (Gendron-Marsolais et al., 2020, 2021) and X-ray cavities linked to the radio lobes of NGC 1275 (3C84) (Fabian et al., 2000). Similarly, the buoyant bubbles ascending around M87 (Churazov et al., 2001) suggest that active galactic nucleus (AGN) feedback plays a substantial role in Virgo. Moreover, the RMH in M87 is remarkable as one of the earliest discovered radio sources (Bolton & Stanley, 1948). Effelsberg observations uncovered a faint radio halo around M86 in Virgo, indicating a past interaction between M86’s intracluster medium (ICM) and low-density gas at the cluster’s periphery (Vollmer et al., 2004).
@@
In these galaxy clusters, point sources contribute to the $\gamma$-ray emission. During the analysis, it is crucial to distinguish this emission from the extended $\gamma$-ray emission. Fermi observations indicate the presence of three high-energy point sources within the Coma galaxy cluster, which may be associated with radio galaxies or starburst galaxies (Baghmanyan et al., 2022). Other studies have detected $\gamma$-ray emissions from NGC 1275 (Aleksić et al., 2012) and IC 310 (Neronov et al., 2010) in the Perseus cluster, as well as from M87 (Aharonian et al., 2003) in the Virgo cluster. Significantly, several recent publications and alerts from LHAASO have reported detections of M87 (LHAASO Collaboration, 2024a), NGC 1275 (LHAASO Collaboration, 2024b), and IC 310 (Xiang et al., 2024a, b). Thus, we incorporate these sources as point sources in our analysis.
@@
In recent years, multiple detectors have observed and placed constraints on the diffuse $\gamma$-ray emission from these clusters. HESS observations of the Coma cluster have constrained $X_{\text{CRp}}$ to values below 0.2, thereby ruling out the most favorable theoretical models (HESS Collaboration & Aharonian, 2009). Fermi observations have reported the detection of diffuse $\gamma$-ray emission from the Coma cluster, estimating $X_{\text{CRp}}$ to be around 1% (Xi et al., 2018). Subsequent observations have confirmed this finding and provided further detailed analyses (Adam et al., 2021; Baghmanyan et al., 2022). However, these observations face challenges in differentiating between multiple point sources and extended sources, and the spectral indices obtained are relatively soft. Observations at other wavelengths, such as radio observations of the synchrotron emission from secondary pairs in the Coma cluster, yield constraints on $X_{\text{CRp}}$ that range from less than 0.01% to 28%, depending on the assumptions made regarding the magnetic field. For the Perseus cluster, MAGIC telescope observations have constrained $X_{\text{CRp}}$ to less than approximately 1-2% (MAGIC Collaboration et al., 2010, 2016, 2012). Preliminary simulations based on the expected performance of CTA indicate that the constraints on $X_{\text{CRp}}$ could be an order of magnitude stronger than those from previous studies (CTA Consortium et al., 2023). HESS observations of the Virgo cluster have also constrained $X_{\text{CRp}}$ to values below 0.32 (HESS Collaboration et al., 2023). These constraints are predicated on assumptions about the spatial and spectral distribution of cosmic rays and are applicable to different spatial scales.
@@
3 Data analysis
3.1 Analysis method
@@
After data selection, the number of events used in this analysis is $4.69\times 10^{9}$ for WCDA and $1.90\times 10^{9}$ for KM2A. For more details regarding the array and the reconstruction of WCDA and KM2A, refer to (Aharonian et al., 2021a, b).
Figure 1: The upper panel shows the WCDA skymaps (1-25 TeV) with the contribution of known point sources subtracted, while the lower panel displays the KM2A skymaps ($>$ 25 TeV). From left to right, they correspond to the Coma, Perseus, and Virgo clusters. The $R_{500}$ radius of these clusters is respectively indicated by gray dashed lines. The known $\gamma$-rays point sources NGC 1275 and IC 310 in the Perseus cluster are respectively marked with cyan squares and upward triangles, while M87 in the Virgo cluster is indicated by a green star.
The WCDA leverages the number of triggered PMT units per event, denoted as $N_{\text{hit}}$, to approximate the event energy. All events are partitioned into six $N_{\text{hit}}$ bins: 60 - 100, 100 - 200, 200 - 300, 300 - 500, 500 - 800, and $>$ 800.
@@
Based on the reconstructed direction of each event, all events are binned into skymaps with a pixel size of $0.1^{\circ}\times 0.1^{\circ}$. The cosmic-ray background is estimated using the direct integration method (Fleysher et al., 2004).
The three-dimensional (3D) algorithm was employed to simultaneously fit both the spectrum and morphology of the source. In this study, we compute the source significance using ${\rm TS}=2\log\left({\frac{\mathcal{L}_{1}}{\mathcal{L}_{0}}}\right)$. Here, $\mathcal{L}_{1}$ represents the maximum likelihood under the alternative hypothesis we intend to test, while $\mathcal{L}_{0}$ denotes the maximum likelihood under the null hypothesis. According to Wilks’ Theorem (Wilks, 1938), the TS value follows a chi-squared distribution with degrees of freedom equal to the difference between those of the null and alternative hypotheses.
To determine the significance in the sky map for this work, we assume that the source is a point source. For WCDA in the energy range of 1-25 TeV, we adopt a power-law spectrum with an index of 2.7; for KM2A at energies above 25 TeV, we use an index of 3.5. In each pixel, the flux is the only free parameter. In line with Wilks’ Theorem, we consider ±$\pm\sqrt{\rm TS}$ as the measure of significance.
@@
3.2 Results
Figure 2: Left panel: The 95% C.L. flux upper limits for the extended emission within the $R_{500}$ radius of the three galaxy clusters, after correction for the EBL, are presented. The darker-blue and lighter-blue arrows represent the data from WCDA and KM2A, respectively. The gray solid and dashed lines represent the $\gamma$-ray fluxes scaled to the strongest constraint for the IC process of secondary electrons produced by the $p-\gamma$ interactions of accelerated cosmic rays with linear and nonlinear DSA scenarios, respectively, as shown in Figs. 7 and 13 of Vannoni et al. (2011). The blue dashed line represents the $\gamma$-ray flux scaled to the strongest constraint for hadronic processes. From top to bottom, the corresponding clusters are the Coma, Perseus, and Virgo clusters. Right panel: The 95% C.L. upper-limit constraints on $X_{\text{CRp}}$ for these three galaxy clusters, under the hypothesis of different CRp spectral indices, are shown. The sequence of galaxy clusters from top to bottom is consistent with the order presented in the left panel. The solid squares in the first plot on the right panel represent the strongest constraints (99% C.L.) previously obtained by HESS (HESS Collaboration & Aharonian, 2009). The hollow pentagrams and crosses represent VERITAS’s and Fermi’s strongest constraints (99% C.L.) on $X_{\text{CRp}}$ within the virial radius of Coma (Arlen et al., 2012). The hollow circles in the second plot on the right denote the 95% C.L. constraints on $X_{\text{CRp}}$ within the virial radius of the Perseus cluster, as determined by MAGIC (MAGIC Collaboration et al., 2016). The hollow squares in the bottom-most plot represent the 99.7% C.L. constraints on $X_{\text{CRp}}$ within the Virgo cluster at $\sim 20\,\text{kpc}$, as determined by HESS (HESS Collaboration et al., 2023). In the bottom two panels, the open crosses represent the 95% C.L. upper limits provided by Fermi (Ackermann et al., 2010).

### v2

@@ Abstract
Galaxy clusters act as reservoirs of high-energy cosmic rays (CRs). As CRs propagate through the intracluster medium, they generate diffuse $\gamma$-rays detectable by arrays such as LHAASO. These $\gamma$-rays result from proton-proton ($pp$) collisions of very high-energy cosmic rays (VHECRs) or inverse Compton (IC) scattering of positron-electron pairs created by $p\gamma$ interactions of ultra-high-energy cosmic rays (UHECRs). We analyzed diffuse $\gamma$-ray emission from the Coma, Perseus, and Virgo clusters using LHAASO data. Diffuse emission was modeled as a disk of radius $R_{500}$ for each cluster while accounting for point sources. No significant diffuse emission was detected, yielding 95% confidence level (C.L.) upper limits on the $\gamma$-ray flux: for WCDA (1–25 TeV) and KM2A (>>25 TeV), less than (49.4, 13.7, 54.0) and (1.34, 1.14, 0.40) $\times 10^{-14}$ ph cm-2 s-1 for Coma, Perseus, and Virgo, respectively. The $\gamma$-ray upper limits can be used to derive model-independent constraints on the integral energy of CRp above 10 TeV (corresponding to the LHAASO observational range $1\,\mathrm{TeV}$ under the $pp$ scenario) to be less than $(1.96,0.59,0.08)\times 10^{61}\,\mathrm{erg}$. The absence of detectable annuli/ring-like structures, indicative of cluster accretion or merging shocks, imposes further constraints on models in which the UHECRs are accelerated in the merging shocks of galaxy clusters.
@@ I Introduction
Galaxy clusters, recognized as the largest gravitationally bound entities in the universe, serve as vital cosmological laboratories for investigating a wide range of astronomical phenomena. These clusters typically encompass hundreds to thousands of galaxies, all enveloped in hot, diffuse intracluster gas that extends across several megaparsecs. Radio observations have uncovered non-thermal radio halos and relics within these clusters, indicating the presence of a population of relativistic particles (ferettiClustersGalaxiesObservational2012).
@@
Cosmic rays within galaxy clusters are generated and accelerated by multiple astrophysical sources. These include supernova-driven galactic winds, active galactic nuclei (AGNs) (hintonRayEmissionAssociated2007), large-scale intergalactic shocks arising from accretion and merger processes (colafrancescoClustersGalaxiesDiffuse1998; ryuCosmologicalShockWaves2003; vannoniAccelerationRadiationUltrahigh2011; brunettiCosmicRaysGalaxy2014), and turbulent intracluster magnetic fields (brunettiAlfvenicReaccelerationRelativistic2005). The magnetic fields in galaxy clusters, typically on the order of a few $\mu G$, effectively confine cosmic rays, allowing their accumulation (vlkNonthermalEnergyContent1996; berezinskyClustersGalaxiesStorage1997a; govoniMAGNETICFIELDSCLUSTERS2004). When these relativistic particles interact with the ambient matter, they can generate detectable high-energy $\gamma$-rays.
@@
Studying galaxy clusters in the $\gamma$-ray band provides crucial insights into the efficiency of particle acceleration, magnetic confinement, and radiation processes involving cosmic rays. Several radiation mechanisms can give rise to $\gamma$-rays in galaxy clusters. The hadronic scenario, which involves inelastic collisions between cosmic ray protons (CRp) and thermal nuclei (dennisonFormationRadioHalos1980), is a major potential $\gamma$-ray radiation mechanism in galaxy clusters. Leptonic scenarios can also generate $\gamma$-rays. Relativistic CRp interact with target photon fields, such as the cosmic microwave background (CMB) and infrared background, producing high-energy $\gamma$-rays (atoyanImplicationsNonthermalOrigin2000). $\gamma$-ray emission can also arise from IC scattering of secondary electrons, which are produced by Bethe-Heitler processes involving relativistic protons and the CMB photon field, if UHECRs are accelerated within galaxy clusters, as predicted by some models (kelnerEnergySpectraGamma2008). Another mechanism involves IC radiation emitted by relativistic cosmic ray electrons (CRe), which originate from interactions between very high-energy (VHE) $\gamma$-rays and the diffuse extragalactic background radiation field.
@@
Previous observations of $\gamma$-rays in the MeV to GeV energy range by EGRET (reimerEGRETUpperLimits2003) provided only upper limits for several galaxy clusters. Similarly, the Whipple telescope reported upper limits at TeV energies for the Perseus and Abell 2029 clusters (perkinsTeVGammaRay2006). More recently, stereoscopic instruments such as H.E.S.S. and VERITAS have also set upper limits from TeV observations (collaborationConstrainingCosmicrayPressure2023c; thehesscollaborationConstraintsMultiTeVParticle2009; arlenCONSTRAINTSCOSMICRAYS2012; perkinsTeVGammaRay2006). The analysis of Fermi LAT data by xiDetectionGammarayEmission2018 revealed extended GeV emission from the direction of the Coma cluster, which was later confirmed by baghmanyanDetailedStudyExtended2022.
@@
In this Letter, we report the LHAASO observations of three nearby galaxy clusters: Coma, Perseus, and Virgo. Leveraging the $\gamma$-ray upper limits obtained from LHAASO observations, we impose stringent constraints on the energy budget of high-energy particles within these clusters above 10 TeV, corresponding to the $\gamma$-ray energy range observed by LHAASO ($1\,\mathrm{TeV}$) under the $pp$ interaction scenario, in a model-independent manner.
@@
The structure of this paper is as follows. In Sec.II, we present the fundamental background information on the selected clusters. In Sec.III, we conduct an in-depth analysis of the LHAASO-WCDA and LHAASO-KM2A data. Finally, in Sec.IV, we discuss the derived cosmic-ray constraints and their astrophysical implications.
II Facts of the galaxy clusters
@@
Coma, Perseus, and Virgo are the three closest massive galaxy clusters, with redshifts of approximately 0.023 (giovanniniHaloRadioSource1993), 0.017 (hitomicollaborationAtmosphericGasDynamics2018), and 0.004 (meiACSVirgoCluster2007), and total masses $M_{500}$ of about $6.13\times 10^{14}M_{\odot}$ (planckcollaborationPlanckIntermediateResults2013), $5.77\times 10^{14}M_{\odot}$ (consortiumProspectsGammaRay2023), and $0.83\times 10^{14}M_{\odot}$ (simionescuWitnessingGrowthNearest2017), respectively (as shown in Tab.A.1). Their proximity and large mass render these clusters ideal targets for studying cosmic rays and dark matter.
@@
Observations in other wavelengths have revealed that all three clusters exhibit radio halos, which are a clear indication of a substantial population of relativistic particles. Specifically, the Coma cluster possesses a giant radio halo, which is typically associated with cluster mergers (cassanoConnectionGiantRadio2010). This halo extends from the core of the Coma cluster and features diffuse radio relics near its virial radius (thierbachDiffuseRadioEmission2003). X-ray observations have shown that NGC 4839 is a subgroup in the south-west region of the Coma cluster, connected to its core by a faint X-ray bridge, suggesting a recent merger event (churazovTempestuousLifeR5002021; malavasiSpiderItsWeb2020). The presence of the giant radio halo, in tandem with the merger activities, implies that shock wave dissipation and turbulence resulting from the merging process play significant roles in injecting non-thermal energy into the system (vannoniAccelerationRadiationUltrahigh2011).
@@
Perseus and Virgo are well-known as relaxed cool-core (CC) clusters (petersonXraySpectroscopyCooling2006), characterized by their dense cores. Perseus showcases a radio mini-halo (RMH) (gendron-marsolaisHighresolutionVLALow2020; gendron-marsolaisVLAResolvesUnexpected2021) and X-ray cavities linked to the radio lobes of NGC 1275 (3C84) (fabianChandraImagingComplex2000). Similarly, the buoyant bubbles ascending around M87 (churazovEvolutionBuoyantBubbles2001) suggest that active galactic nucleus (AGN) feedback plays a substantial role in Virgo. Moreover, the RMH in M87 is remarkable as one of the earliest discovered radio sources (boltonVariableSourceRadio1948). Effelsberg observations uncovered a faint radio halo around M86 in Virgo, indicating a past interaction between M86’s intracluster medium (ICM) and low-density gas at the cluster’s periphery (vollmerDetectionRadioHalo2004).
@@
In these galaxy clusters, point sources contribute to the $\gamma$-ray emission. During the analysis, it is crucial to distinguish this emission from the extended $\gamma$-ray emission. Fermi observations indicate the presence of three high-energy point sources within the Coma galaxy cluster, which may be associated with radio galaxies or starburst galaxies (baghmanyanDetailedStudyExtended2022). Other studies have detected $\gamma$-ray emissions from NGC 1275 (aleksicDetectionVeryhighEnergy2012) and IC 310 (neronovVeryHighenergyRay2010) in the Perseus cluster, as well as from M87 (aharonianGiantRadioGalaxy2003a) in the Virgo cluster. Significantly, several recent publications and alerts from LHAASO have reported detections of M87 (collaborationDetectionVeryHighenergy2024), NGC 1275 (wangDetectionTwoTeV2024), and IC 310 (xiangLHAASODetectsRapid2024a; xiangLHAASODetectionRenewed2024). Thus, we incorporate these sources as point sources in our analysis.
@@
In recent years, multiple detectors have observed and placed constraints on the diffuse $\gamma$-ray emission from these clusters. HESS observations of the Coma cluster have constrained $X_{\text{CRp}}$ to values below 0.2, thereby ruling out the most favorable theoretical models (thehesscollaborationConstraintsMultiTeVParticle2009). Fermi observations have reported the detection of diffuse $\gamma$-ray emission from the Coma cluster, estimating $X_{\text{CRp}}$ to be around 1% (xiDetectionGammarayEmission2018). Subsequent observations have confirmed this finding and provided further detailed analyses (adamRayDetectionComa2021; baghmanyanDetailedStudyExtended2022). However, these observations face challenges in differentiating between multiple point sources and extended sources, and the spectral indices obtained are relatively soft. Observations at other wavelengths, such as radio observations of the synchrotron emission from secondary pairs in the Coma cluster, yield constraints on $X_{\text{CRp}}$ that range from less than 0.01% to 28%, depending on the assumptions made regarding the magnetic field. For the Perseus cluster, MAGIC telescope observations have constrained $X_{\text{CRp}}$ to less than approximately 1-2% (themagiccollaborationMAGICGammarayTelescope2010; magiccollaborationDeepObservationNGC2016; magiccollaborationConstrainingCosmicRays2012). Preliminary simulations based on the expected performance of CTA indicate that the constraints on $X_{\text{CRp}}$ could be an order of magnitude stronger than those from previous studies (consortiumProspectsGammaRay2023). HESS observations of the Virgo cluster have also constrained $X_{\text{CRp}}$ to values below 0.32 (collaborationConstrainingCosmicrayPressure2023c). These constraints are predicated on assumptions about the spatial and spectral distribution of cosmic rays and are applicable to different spatial scales.
@@
III Data analysis
III.1 Analysis method
@@
After data selection, the number of events used in this analysis is $4.69\times 10^{9}$ for WCDA and $1.90\times 10^{9}$ for KM2A. For more details regarding the array and the reconstruction of WCDA and KM2A, refer to (aharonianObservationCrabNebula2021a; aharonianPerformanceLHAASOWCDAObservation2021b).
Figure 1: The upper panel shows the WCDA skymaps (1-25 TeV) with the contribution of known point sources subtracted, while the lower panel displays the KM2A skymaps (>> 25 TeV). From left to right, they correspond to the Coma, Perseus, and Virgo clusters. The $R_{500}$ radius of these clusters is respectively indicated by gray dashed lines. The known $\gamma$-rays point sources NGC 1275 and IC 310 in the Perseus cluster are respectively marked with cyan squares and upward triangles, while M87 in the Virgo cluster is indicated by a green star.
The WCDA leverages the number of triggered PMT units per event, denoted as $N_{\text{hit}}$, to approximate the event energy. All events are partitioned into six $N_{\text{hit}}$ bins: 60 - 100, 100 - 200, 200 - 300, 300 - 500, 500 - 800, and >> 800.
@@
Based on the reconstructed direction of each event, all events are binned into skymaps with a pixel size of $0.1^{\circ}\times 0.1^{\circ}$. The cosmic-ray background is estimated using the direct integration method (fleysherTestsStatisticalSignificance2004b).
The three-dimensional (3D) algorithm was employed to simultaneously fit both the spectrum and morphology of the source. In this study, we compute the source significance using ${\rm TS}=2\log\left({\frac{\mathcal{L}_{1}}{\mathcal{L}_{0}}}\right)$. Here, $\mathcal{L}_{1}$ represents the maximum likelihood under the alternative hypothesis we intend to test, while $\mathcal{L}_{0}$ denotes the maximum likelihood under the null hypothesis. According to Wilks’ Theorem (wilksLargeSampleDistributionLikelihood1938a), the TS value follows a chi-squared distribution with degrees of freedom equal to the difference between those of the null and alternative hypotheses.
To determine the significance in the sky map for this work, we assume that the source is a point source. For WCDA in the energy range of 1-25 TeV, we adopt a power-law spectrum with an index of 2.7; for KM2A at energies above 25 TeV, we use an index of 3.5. In each pixel, the flux is the only free parameter. In line with Wilks’ Theorem, we consider $\pm\sqrt{\rm TS}$ as the measure of significance.
@@
III.2 Results
Figure 2: Left panel: The 95% C.L. flux upper limits for the extended emission within the $R_{500}$ radius of the three galaxy clusters, after correction for the EBL, are presented. The darker-blue and lighter-blue arrows represent the data from WCDA and KM2A, respectively. The gray solid and dashed lines represent the $\gamma$-ray fluxes scaled to the strongest constraint for the IC process of secondary electrons produced by the $p-\gamma$ interactions of accelerated cosmic rays with linear and nonlinear DSA scenarios, respectively, as shown in Figs. 7 and 13 of vannoniAccelerationRadiationUltrahigh2011. The blue dashed line represents the $\gamma$-ray flux scaled to the strongest constraint for hadronic processes. From top to bottom, the corresponding clusters are the Coma, Perseus, and Virgo clusters. Right panel: The 95% C.L. upper-limit constraints on $X_{\text{CRp}}$ for these three galaxy clusters, under the hypothesis of different CRp spectral indices, are shown. The sequence of galaxy clusters from top to bottom is consistent with the order presented in the left panel. The solid squares in the first plot on the right panel represent the strongest constraints (99% C.L.) previously obtained by HESS (thehesscollaborationConstraintsMultiTeVParticle2009). The hollow pentagrams and crosses represent VERITAS’s and Fermi’s strongest constraints (99% C.L.) on $X_{\text{CRp}}$ within the virial radius of Coma (arlenCONSTRAINTSCOSMICRAYS2012). The hollow circles in the second plot on the right denote the 95% C.L. constraints on $X_{\text{CRp}}$ within the virial radius of the Perseus cluster, as determined by MAGIC (magiccollaborationDeepObservationNGC2016). The hollow squares in the bottom-most plot represent the 99.7% C.L. constraints on $X_{\text{CRp}}$ within the Virgo cluster at $\sim 20\,\text{kpc}$, as determined by HESS (collaborationConstrainingCosmicrayPressure2023c). In the bottom two panels, the open crosses represent the 95% C.L. upper limits provided by Fermi (ackermannGeVGAMMARAYFLUX2010).